OpenStudy (anonymous):

The pyramids at Giza include several square-based pyramids that were built in honor of the Egyptian gods. One of these incredible monuments is the Khafre pyramid, built as part of a temple complex, with the famous Sphinx at its base. The base length of this pyramid is feet, and its height is feet. Massive dimensions indeed! Calculate the volume of the Khafre pyramid. Show the steps of your solution and explain your work. The pyramids at Giza include several square-based pyramids that were built in honor of the Egyptian gods. One of these incredible monuments is the Khafre pyramid, built as part of a temple complex, with the famous Sphinx at its base. The base length of this pyramid is feet, and its height is feet. Massive dimensions indeed! Calculate the volume of the Khafre pyramid. Show the steps of your solution and explain your work. @Mathematics

OpenStudy (anonymous):

Please help!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

The volume of a square-based pyramid is found by the formula: \[\Large V = \frac{1}{3}{b^2}h\] Since you haven't included the base or the height there's not much more I can help with.

OpenStudy (anonymous):

705 ft and 471 ft

OpenStudy (anonymous):

Then plug it into the formula.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The pyramids at Giza include several square-based pyramids that were built in honor of the Egyptian gods. One of these incredible monuments is the Khafre pyramid, built as part of a temple complex, with the famous Sphinx at its base. The base length of this pyramid is 705 feet, and its height is 471 feet. Massive dimensions indeed! Calculate the volume of the Khafre pyramid. Show the steps of your solution and explain your work

OpenStudy (anonymous):

Yes, so use the formula... =) \[\Large \begin{array}{l} V = \frac{1}{3}{705^2} \cdot 471\\ = \frac{1}{3}497025 \cdot 471\\ = \frac{1}{3}234098775\\ = 78032925\\ = 7.8 \times {10^7}{\rm{ f}}{{\rm{t}}^3} \end{array}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No worries.